OpenStudy (anonymous):

starting with V=Voe^-t/CR derive t1/2=CRln2

OpenStudy (irishboy123):

V/Vo = e ^ (-t/CR) -t/CR = ln (V/Vo) t = - CR ln (V/Vo) = CR ln (Vo/V) if V = 1/2 * Vo, then Vo = 2V, so Vo/V = 2, ERGO t(1/2) = CR ln(2) t(1/2) is half life, time for mass to decay exponentially to half it's original amount.

OpenStudy (anonymous):

sorry i left i had to fix something

OpenStudy (anonymous):

hope this helped bye